Two bodies of masses 10,000kgs and 1000kg respectively is seperated by 1m. At what distance from the smaller body, the intensity of gravitational field will be zero

Grade:11

1 Answers

onkar kumar
11 years ago

Let x be the distance from smaller object where field is zero.

G(10000)/(1-x)^2=G(1000)/x^2

->  sqrt(10)/(1-x)=1/x

-> x= 1/(sqrt(10)+1)

      = (sqrt(10)-1)/9

      =0.24m